# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ IV.4 Exclusive Z boson decays into two mesons
 $2.3$$\times~{}10^{-14}$ NRQCD+LDME [165]
 $2.7$$\times~{}10^{-11}$ NRQCD [159]
@@
 $8.3$$\times~{}10^{-14}$ NRQCD+LDME [165]
 $(4.7\pm 2.0)$$\times~{}10^{-13}$ LC+LCDA [165]
[…]
 $3.5$$\times~{}10^{-15}$ NRQCD+LDME [165]
$\rm h_{c}$  $1.5$$\times~{}10^{-12}$ NRQCD+LDME [165]
 $(9.5\pm 5.0)$$\times~{}10^{-12}$ LC+LCDA [165]
[…]
 $1.4$$\times~{}10^{-13}$ NRQCD+LDME [165]
 $(9.3\pm 4.0)$$\times~{}10^{-13}$ LC+LCDA [165]
$\rm\chi_{c0}$ + $\rm\chi_{c1}$  $7.6$$\times~{}10^{-14}$ NRQCD+LDME [165]
 $(1.4\pm 1.0)$$\times~{}10^{-12}$ LC+LCDA [165]
$\rm h_{c}$  $3.5$$\times~{}10^{-16}$ NRQCD+LDME [165]
$\rm\chi_{c2}$  $6.4$$\times~{}10^{-15}$ NRQCD+LDME [165]
$\rm\chi_{c1}$ + $\rm\chi_{c1}$  $3.9$$\times~{}10^{-16}$ NRQCD+LDME [165]
$\rm h_{c}$  $2.9$$\times~{}10^{-14}$ NRQCD+LDME [165]
 $(6.1\pm 5.0)$$\times~{}10^{-13}$ LC+LCDA [165]
$\rm\chi_{c2}$  $1.3$$\times~{}10^{-13}$ NRQCD+LDME [165]
 $(2.8\pm 2.0)$$\times~{}10^{-12}$ LC+LCDA [165]
$\rm h_{c}$ + $\rm h_{c}$  $9.9$$\times~{}10^{-17}$ NRQCD+LDME [165]
$\rm\chi_{c2}$  $2.3$$\times~{}10^{-16}$ NRQCD+LDME [165]
$\rm\chi_{c2}$ + $\rm\chi_{c2}$  $1.3$$\times~{}10^{-16}$ NRQCD+LDME [165]
Figure 23: Branching fractions (in negative log scale) of exclusive Z boson decays into two light mesons (upper) or two charmonium mesons (lower). The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.
Table 20: Compilation of exclusive Z decays to one bottomonium meson plus another meson (or anything else, in a few cases). For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last column indicates whether the decay can be produced at FCC-ee.
@@
$Z\,\to$ M $+$ M Branching fraction Framework 2024 HL-LHC at FCC-ee
$\rm Z\,\to$ $\rm\Upsilon(1S)$ + $\rm J/\psi$  $4.6$$\times~{}10^{-11}$ NRQCD [159]
$\rm\eta_{b}(1S)$  $(1.9\pm 0.2)$$\times~{}10^{-11}$ NRQCD/NRCSM [160]
$\rm\Upsilon(1S)$  $6.2$$\times~{}10^{-11}$ NRQCD(NLO)+LDME [168] $<$ $1.8$×$\times~{}10^{-6}$ [107] $\lesssim 2.7\times~{}10^{-7}$
 $\left(4.4^{+0.6}_{-0.3}\right)$$\times~{}10^{-13}$ NRQCD/NRCSM [160]
$\rm\Upsilon(2S)$  $5.6$$\times~{}10^{-11}$ NRQCD(NLO)+LDME [168]
$\rm\Upsilon(3S)$  $4.1$$\times~{}10^{-11}$ NRQCD(NLO)+LDME [168]
$\rm\Upsilon(2S)$ + $\rm\Upsilon(2S)$  $1.4$$\times~{}10^{-11}$ NRQCD(NLO)+LDME [168]
$\rm\Upsilon(3S)$  $1.9$$\times~{}10^{-11}$ NRQCD(NLO)+LDME [168]
$\rm\Upsilon(3S)$ + $\rm\Upsilon(3S)$  $6.8$$\times~{}10^{-12}$ NRQCD(NLO)+LDME [168]
$\rm\Upsilon(mS)$ + $\rm\Upsilon(nS)$  $2.1$$\times~{}10^{-12}$ NRQCD [159] $<$ $3.9$×$\times~{}10^{-7}$ [107] $\lesssim 5.9\times~{}10^{-8}$
@@
Figure 24: Branching fractions (in negative log scale) of exclusive $\rm Z\to b$-meson+\,+\,+X decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.
V Rare W boson decays
V.1 Exclusive W decays into a photon plus a meson
Figure 25 displays the two generic diagrams contributing to exclusive decays of a W boson into a photon (the only gauge boson kinematically accessible for the least massive weak boson) plus a charged meson. The partial widths for these decays have been evaluated within the SCET framework in Ref. [44], and can be written as
Re $\text{Re }F_{1}^{\mathrm{M}}$ $=0.94-1.65\ a_{1}^{\mathrm{M}}(1\text{ GeV})+0.42\ a_{2}^{\mathrm{M}}(1\text{ GeV})+\dots$
Re $\text{Re }F_{2}^{\mathrm{M}}$ $=0.85-0.55\ a_{1}^{\mathrm{M}}(1\text{ GeV})+1.25\ a_{2}^{\mathrm{M}}(1\text{ GeV})+\dots,$ (49)
with $a_{i}^{\mathrm{M}}(\mu)$ being the $i$th Gegenbauer moment of the meson $\rm M$, evaluated at the renormalization scale $\mu$. Table 21 compiles the corresponding theoretical predictions and experimental limits for branching ratios of the eight radiative exclusive decay modes considered in Ref. [44], to which we have added the result for the $\rm B_{c}^{\pm}$ meson. The Gegenbauer moments for this latter meson, with one-loop QCD radiative corrections and relativistic corrections [173], are: $a_{1}^{\mathrm{B_{c}}}(\mu_{\rm B_{c}})=0.32,\ a_{2}^{\mathrm{B_{c}}}(\mu_{\rm B_{c}})=-0.34$ at scales $\mu_{\rm B_{c}}=5.6\text{ GeV}$. These values can be evolved to lower scales using the LO renormalization group equation for $N_{\mathrm{f}}=5$ [44], yielding $a_{1}^{\rm B_{c}}(1\text{ GeV})=1.35\ a_{2}^{\rm B_{c}}(\mu_{\rm B_{c}}),\ a_{2}^{\rm B_{c}}(1\text{ GeV})=1.6\ a_{2}^{\rm B_{c}}(\mu_{\rm B_{c}})$, for $\alpha_{\mathrm{s}}(\mu_{\rm B_{c}})=0.21,\ \alpha_{\mathrm{s}}(1\text{ GeV})=0.4$, which we use in Eqs. (48)-(49) to obtain the result quoted in the last row of Table 21. The theoretical branching fractions are in the $\mathcal{O}(10^{-8}$$10^{-13})$ range, whereas current experimental upper bounds are three to five orders-of-magnitude larger, $\mathcal{O}(10^{-4}$$10^{-6})$. Four channels have been searched for to date at the LHC (3 by CMS, 1 by ATLAS, and 1 by LHCb), and the HL-LHC will improve the existing limits by at least one order-of-magnitude, but still far from any possible observation. The FCC-ee will be able to produce events for three decay modes, and the FCC-hh will produce all of them. The results listed in Table 21 are presented in graphical form in Fig. 26.
Table 21: Compilation of exclusive W boson decays to a photon plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
Exp. limits Producible at
$\rm W^{\pm}\,\to\,\gamma$ + $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm W^{\pm}\,\to\,\gamma$ + $\rm\pi^{\pm}$  $(4.0\pm 0.8)$$\times~{}10^{-9}$ SCET+LCDA [44] $<$ $1.9$×$\times~{}10^{-6}$ [174, 175] $\lesssim 2.9\times~{}10^{-7}$
$\rm\rho^{\pm}$  $(8.74\pm 1.91)$$\times~{}10^{-9}$ SCET+LCDA [44] $<$ $5.2$×$\times~{}10^{-6}$ [174] $\lesssim 7.9\times~{}10^{-7}$
$\rm K^{\pm}$  $(3.25\pm 0.69)$$\times~{}10^{-10}$ SCET+LCDA [44] $<$ $1.7$×$\times~{}10^{-6}$ [174] $\lesssim 2.6\times~{}10^{-7}$
$\rm K^{*\pm}$  $(4.78\pm 1.15)$$\times~{}10^{-10}$ SCET+LCDA [44]
$\rm D^{\pm}$  $\left(1.38^{+0.51}_{-0.33}\right)$$\times~{}10^{-9}$ SCET+LCDA [44]
$\rm D_{s}^{\pm}$  $4.7$$\times~{}10^{-9}$ NRQCD+LDME [176] $<$ $6.5$×$\times~{}10^{-4}$ [144] $\lesssim 1.2\times~{}10^{-5}$
 $3.4$$\times~{}10^{-9}$ LC+LCDA [176]
 $\left(3.66^{+1.49}_{-0.85}\right)$$\times~{}10^{-8}$ SCET+LCDA [44]
$\rm D_{s}^{*\pm}$  $8.9$$\times~{}10^{-10}$ NRQCD+LDME [176]
 $3.4$$\times~{}10^{-9}$ LC+LCDA [176]
$\rm B^{\pm}$  $\left(2.616^{+3.146}_{-1.330}\right)$$\times~{}10^{-12}$ HQET+LCDA [177]
 $\left(1.99^{+2.49}_{-0.82}\right)$$\times~{}10^{-12}$ SCET+LCDA111111updated inputs from [177] for the result from [44] [177]
 $\left(1.55^{+0.79}_{-0.60}\right)$$\times~{}10^{-12}$ SCET+LCDA [44]
@@ V.2 Exclusive W decays into two mesons
[…]
@@
Table 22: Compilation of exclusive W boson decays into two mesons. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm W^{\pm}\,\to$ M + M Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm W^{\pm}\,\to$ $\rm\eta_{c}$ + $\rm D_{s}^{\pm}$  $2.1$$\times~{}10^{-12}$ NRCSM+LCDA [179]
 $\left(1.31^{+0.34}_{-0.22}\right)$$\times~{}10^{-11}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $3.0$$\times~{}10^{-12}$ NRCSM+LCDA [179]
 $\left(1.48^{+0.36}_{-0.22}\right)$$\times~{}10^{-11}$ LC+LCDA [179]
$\rm J/\psi$ + $\rm D_{s}^{\pm}$  $2.6$$\times~{}10^{-12}$ NRQCD+LDME [176]
 $2.1$$\times~{}10^{-12}$ NRCSM+LCDA [179]
 $\left(1.8^{+0.4}_{-0.2}\right)$$\times~{}10^{-11}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $1.7$$\times~{}10^{-12}$ NRQCD+LDME [176]
 $3.0$$\times~{}10^{-12}$ NRCSM+LCDA [179]
 $\left(2.03^{+0.46}_{-0.24}\right)$$\times~{}10^{-11}$ LC+LCDA [179]
$\rm\psi(2S)$ + $\rm D_{s}^{\pm}$  $5.1$$\times~{}10^{-11}$ NRQCD+LDME [176]
$\rm D_{s}^{*\pm}$  $7.4$$\times~{}10^{-12}$ NRQCD+LDME [176]
$\rm\chi_{c0}$ + $\rm D_{s}^{\pm}$  $9.4$$\times~{}10^{-14}$ NRQCD+LDME [176]
 $4.7$$\times~{}10^{-14}$ NRCSM+LCDA [179]
 $\left(7.1^{+3.5}_{-3.1}\right)$$\times~{}10^{-13}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $1.2$$\times~{}10^{-13}$ NRQCD+LDME [176]
 $8.1$$\times~{}10^{-14}$ NRCSM+LCDA [179]
 $\left(8.0^{+3.7}_{-3.1}\right)$$\times~{}10^{-13}$ LC+LCDA [179]
$\rm\chi_{c1}$ + $\rm D_{s}^{\pm}$  $2.0$$\times~{}10^{-13}$ NRQCD+LDME [176]
 $2.9$$\times~{}10^{-13}$ NRCSM+LCDA [179]
 $\left(7.83^{+3.40}_{-3.05}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $2.0$$\times~{}10^{-13}$ NRQCD+LDME [176]
 $4.0$$\times~{}10^{-13}$ NRCSM+LCDA [179]
 $\left(8.83^{+3.50}_{-3.06}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm h_{c}$ + $\rm D_{s}^{\pm}$  $1.4$$\times~{}10^{-13}$ NRCSM+LCDA [179]
 $\left(2.13^{+0.97}_{-0.83}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $2.0$$\times~{}10^{-13}$ NRCSM+LCDA [179]
 $\left(2.4^{+1.1}_{-0.9}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm\chi_{c2}$ + $\rm D_{s}^{\pm}$  $3.9$$\times~{}10^{-14}$ NRQCD+LDME [176]
 $9.6$$\times~{}10^{-14}$ NRCSM+LCDA [179]
 $\left(1.42^{+0.62}_{-0.53}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm D_{s}^{*\pm}$  $3.9$$\times~{}10^{-14}$ NRQCD+LDME [176]
 $1.4$$\times~{}10^{-13}$ NRCSM+LCDA [179]
 $\left(1.6^{+0.7}_{-0.6}\right)$$\times~{}10^{-12}$ LC+LCDA [179]
$\rm B_{s}^{0}$ + $\rm B_{c}^{\pm}$  $2.5$$\times~{}10^{-12}$ NRQCD+LDME [176]
$\rm B_{c}^{*\pm}$  $2.0$$\times~{}10^{-12}$ NRQCD+LDME [176]
$\rm B_{s}^{*}$ + $\rm B_{c}^{\pm}$  $2.7$$\times~{}10^{-12}$ NRQCD+LDME [176]
$\rm B_{c}^{*\pm}$  $2.7$$\times~{}10^{-12}$ NRQCD+LDME [176]
$\rm J/\psi$ + $\rm D^{\pm}$  $\mathcal{O}$ $(10^{-13})$ LC+$\delta$ appr. (this work)
$\rm\Upsilon(1S)$ + $\rm B^{\pm}$  $\mathcal{O}$ $(10^{-16})$ LC+$\delta$ appr. (this work)
$\rm B_{c}^{\pm}$  $\mathcal{O}$ $(10^{-14})$ LC+$\delta$ appr. (this work)

### v4

@@ IV.4 Exclusive Z boson decays into two mesons
$\left(4.4^{+0.6}_{-0.3}\right)$ $\times~10^{-13}$ NRQCD/NRCSM [Gao:2022mwa]
$\rm\Upsilon(2S)$ $(5.6\pm 0.2)$ $\times~10^{-11}$ NRQCD(NLO)+LDME [Li:2024zun]
$\rm\Upsilon(3S)$ $(4.1\pm 0.2)$ $\times~10^{-11}$ NRQCD(NLO)+LDME [Li:2024zun]
$\rm\Upsilon(2S)$ + $\rm\Upsilon(2S)$ $(1.4\pm 0.1)$ $\times~10^{-11}$ NRQCD(NLO)+LDME [Li:2024zun]
$\rm\Upsilon(3S)$ $(1.9\pm 0.1)$ $\times~10^{-11}$ NRQCD(NLO)+LDME [Li:2024zun]
$\rm\Upsilon(3S)$ + $\rm\Upsilon(3S)$ $(6.8\pm 0.3)$ $\times~10^{-12}$ NRQCD(NLO)+LDME [Li:2024zun]
@@
Figure 24: Branching fractions (in negative log scale) of exclusive $\rm Z\to\Upsilon(nS)\,+\,$M decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.
V Rare W boson decays
V.1 Exclusive W decays into a photon plus a meson
Figure 25 displays the two generic diagrams contributing to exclusive decays of a W boson into a photon (the only gauge boson kinematically accessible for the least massive weak boson) plus a charged meson. The partial widths for these decays have been evaluated within the SCET framework in Ref. [Grossman:2015cak], and can be written as
$\text{Re }F_{1}^{\mathrm{M}}$ $=0.94-1.65\ a_{1}^{\mathrm{M}}(1\text{ GeV})+0.42\ a_{2}^{\mathrm{M}}(1\text{ GeV})+\dots$
$\text{Re }F_{2}^{\mathrm{M}}$ $=0.85-0.55\ a_{1}^{\mathrm{M}}(1\text{ GeV})+1.25\ a_{2}^{\mathrm{M}}(1\text{ GeV})+\dots,$ (49)
with $a_{i}^{\mathrm{M}}(\mu)$ being the $i$th Gegenbauer moment of the meson $\rm M$, evaluated at the renormalization scale $\mu$. Table 21 compiles the corresponding theoretical predictions and experimental limits for branching ratios of the eight radiative exclusive decay modes considered in Ref. [Grossman:2015cak], to which we have added the result for the $\rm B_{c}^{\pm}$ meson. The Gegenbauer moments for this latter meson, with one-loop QCD radiative corrections and relativistic corrections [Wang:2017bgv], are: $a_{1}^{\mathrm{B_{c}}}(\mu_{\rm B_{c}})=0.32,\ a_{2}^{\mathrm{B_{c}}}(\mu_{\rm B_{c}})=-0.34$ at scales $\mu_{\rm B_{c}}=5.6\text{ GeV}$. These values can be evolved to lower scales using the LO renormalization group equation for $N_{\mathrm{f}}=5$ [Grossman:2015cak], yielding $a_{1}^{\rm B_{c}}(1\text{ GeV})=1.35\ a_{2}^{\rm B_{c}}(\mu_{\rm B_{c}}),\ a_{2}^{\rm B_{c}}(1\text{ GeV})=1.6\ a_{2}^{\rm B_{c}}(\mu_{\rm B_{c}})$, for $\alpha_{\mathrm{s}}(\mu_{\rm B_{c}})=0.21,\ \alpha_{\mathrm{s}}(1\text{ GeV})=0.4$, which we use in Eqs. (48)–(49) to obtain the result quoted in the last row of Table 21. The theoretical branching fractions are in the $\mathcal{O}(10^{-8}$$10^{-13})$ range, whereas current experimental upper bounds are three to five orders-of-magnitude larger, $\mathcal{O}(10^{-4}$$10^{-6})$. Four channels have been searched for to date at the LHC (3 by CMS, 1 by ATLAS, and 1 by LHCb), and the HL-LHC will improve the existing limits by at least one order-of-magnitude, but still far from any possible observation. The FCC-ee will be able to produce events for three decay modes, and the FCC-hh will produce all of them. The results listed in Table 21 are presented in graphical form in Fig. 26.
Table 21: Compilation of exclusive W boson decays to a photon plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@ V.1 Exclusive W decays into a photon plus a meson
$\rm W^{\pm}\,\to\,\gamma$ + $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
@@
$\rm K^{*\pm}$ $(4.78\pm 1.15)$ $\times~10^{-10}$ SCET+LCDA [Grossman:2015cak]
$\rm D^{\pm}$ $\left(1.38^{+0.51}_{-0.33}\right)$ $\times~10^{-9}$ SCET+LCDA [Grossman:2015cak]
@@
$3.4$ $\times~10^{-9}$ LC+LCDA [Bagdatova:2023etj]
$\left(3.66^{+1.49}_{-0.85}\right)$ $\times~10^{-8}$ SCET+LCDA [Grossman:2015cak]
$\rm D_{s}^{*\pm}$ $8.9$ $\times~10^{-10}$ NRQCD+LDME [Bagdatova:2023etj]
$3.4$ $\times~10^{-9}$ LC+LCDA [Bagdatova:2023etj]
$\rm B^{\pm}$ $\left(2.616^{+3.146}_{-1.330}\right)$ $\times~10^{-12}$ HQET+LCDA [Beneke:2023nmj]
$\left(1.99^{+2.49}_{-0.82}\right)$ $\times~10^{-12}$ SCET+LCDA131313updated inputs from [Beneke:2023nmj] for the result from [Grossman:2015cak] [Beneke:2023nmj]
$\left(1.55^{+0.79}_{-0.60}\right)$ $\times~10^{-12}$ SCET+LCDA [Grossman:2015cak]
@@ V.2 Exclusive W decays into two mesons
Figure 27 shows the schematic diagrams of the exclusive decay of the W boson into two mesons proceeding through quark decays (left and center), or through intermediate $\rm W^{*},\,\gamma^{*}\to M$ transitions (right). Table 22 lists the corresponding theoretical predictions and experimental limits for concrete decay modes, all of them involving charm and/or bottom quarks mesons. We are not aware of any study of exclusive W-boson decays involving one or two light mesons in the literature [^14] The exclusive three-pion decay, $\rm W^{\pm}\to\pi^{+}\pi^{-}\pi^{\pm}$, estimated to have a $\mathcal{B}\approx 10^{-7}$ [Melia:2016knk], has also been searched-for by the CMS experiment at the LHC [CMS:2019vaj].. The expected rates are truly tiny, of order $\mathcal{O}(10^{-11}\text{--}10^{-14})$, with relatively large uncertainties. The work of [Luchinsky:2021amp] compared the branching fractions predicted within two approaches (LC and NRCSM), which differ by factors of 5–10, whereas Ref. [Bagdatova:2023etj] computed those within the NRQCD framework. For decays with a small ratio $m_{\mathrm{M}}/m_{\mathrm{W}}$, the results from the LC approach should be preferable. Differences among predictions for the same process appear in some cases as due to not consistently considering all diagrams of Fig. 27. The effect of the center diagram with a virtual photon in Fig. 27 was considered in [Bagdatova:2023etj], and its contribution is equal or larger than the direct (left) diagram in Fig. 27. None of the two theoretical works that have so far studied these processes, has considered the middle diagram of Fig. 27 with the $\rm W^{*}$ boson exchange nor the rightmost diagram of Fig. 27. Using the $\delta$-approximation expression for the decay widths given in Ref. [Luchinsky:2021amp], we can see that the $\rm W\to M1+M2$ decay widths are proportional to $|V_{\rm q_{1}q_{2}}|^{2}\times f^{2}_{\rm M_{1}}\times f^{2}_{\rm M_{2}}$, where $f_{\rm M_{1}}$ and $f_{\rm M_{2}}$ are the decay constants of the two mesons. Using such a dependence, we can estimate the order-of-magnitude of other double-meson decay rates not computed to date. The most probable group of decays is the CKM-favoured $\rm W\to c+s\to\psi(\textit{n}S)+D^{(*)}_{s},B^{(*)}_{c}+B^{(*)}_{s}$, already listed in Table 22. The next most probable group of decays involving quarkonia is $\rm W\to c+d\to(cq)+(qd)$, that leads to the $\rm\mathcal{B}(W^{\pm}\to\mathrm{J}/\psi+D^{\pm})\approx\mathcal{O}(10^{-13})$ estimate. Following that, we have the decays $\rm W\to b+c\to(bq)+(qc)$, that result in $\rm\mathcal{B}(W^{\pm}\to\Upsilon(1S)+B_{c}^{\pm})\approx\mathcal{O}(10^{-14})$. Finally, the least probable decay group involving bottom mesons is the CKM-suppressed $\rm W\to b+u\to(bq)+(qu)$, yielding $\rm\mathcal{B}(W\to\Upsilon(1S)+B)\approx\mathcal{O}(10^{-16})$.
@@
Table 22: Compilation of exclusive W boson decays into two mesons. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm W^{\pm}\,\to$ M + M Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
$\rm W^{\pm}\,\to$ $\rm\eta_{c}$ + $\rm D_{s}^{\pm}$ $2.1$ $\times~10^{-12}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(1.31^{+0.34}_{-0.22}\right)$ $\times~10^{-11}$ LC+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $3.0$ $\times~10^{-12}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(1.48^{+0.36}_{-0.22}\right)$ $\times~10^{-11}$ LC+LCDA [Luchinsky:2021amp]
[…]
$\rm D_{s}^{\pm}$ $2.6$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\left(1.8^{+0.4}_{-0.2}\right)$ $\times~10^{-11}$ LC+LCDA [Luchinsky:2021amp]
$2.1$ $\times~10^{-12}$ NRCSM+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $3.0$ $\times~10^{-12}$ NRCSM+LCDA [Luchinsky:2021amp]
$1.7$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\left(2.03^{+0.46}_{-0.24}\right)$ $\times~10^{-11}$ LC+LCDA [Luchinsky:2021amp]
$\rm\psi(2S)$ + $\rm D_{s}^{\pm}$ $5.1$ $\times~10^{-11}$ NRQCD+LDME [Bagdatova:2023etj]
$\rm D_{s}^{*\pm}$ $7.4$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\rm\chi_{c0}$ + $\rm D_{s}^{\pm}$ $9.4$ $\times~10^{-14}$ NRQCD+LDME [Bagdatova:2023etj]
$4.7$ $\times~10^{-14}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(7.1^{+3.5}_{-3.1}\right)$ $\times~10^{-13}$ LC+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $1.2$ $\times~10^{-13}$ NRQCD+LDME [Bagdatova:2023etj]
$8.1$ $\times~10^{-14}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(8.0^{+3.7}_{-3.1}\right)$ $\times~10^{-13}$ LC+LCDA [Luchinsky:2021amp]
$\rm\chi_{c1}$ + $\rm D_{s}^{\pm}$ $2.0$ $\times~10^{-13}$ NRQCD+LDME [Bagdatova:2023etj]
$2.9$ $\times~10^{-13}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(7.83^{+3.40}_{-3.05}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $2.0$ $\times~10^{-13}$ NRQCD+LDME [Bagdatova:2023etj]
$4.0$ $\times~10^{-13}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(8.83^{+3.50}_{-3.06}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm h_{c}$ + $\rm D_{s}^{\pm}$ $1.4$ $\times~10^{-13}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(2.13^{+0.97}_{-0.83}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $2.0$ $\times~10^{-13}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(2.4^{+1.1}_{-0.9}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm\chi_{c2}$ + $\rm D_{s}^{\pm}$ $3.9$ $\times~10^{-14}$ NRQCD+LDME [Bagdatova:2023etj]
$9.6$ $\times~10^{-14}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(1.42^{+0.62}_{-0.53}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm D_{s}^{*\pm}$ $3.9$ $\times~10^{-14}$ NRQCD+LDME [Bagdatova:2023etj]
$1.4$ $\times~10^{-13}$ NRCSM+LCDA [Luchinsky:2021amp]
$\left(1.6^{+0.7}_{-0.6}\right)$ $\times~10^{-12}$ LC+LCDA [Luchinsky:2021amp]
$\rm B_{s}^{0}$ + $\rm B_{c}^{\pm}$ $2.5$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\rm B_{c}^{*\pm}$ $2.0$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\rm B_{s}^{*}$ + $\rm B_{c}^{\pm}$ $2.7$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
$\rm B_{c}^{*\pm}$ $2.7$ $\times~10^{-12}$ NRQCD+LDME [Bagdatova:2023etj]
[…]